How long would it take a population of 3.5% to double itself
Answer by stanbon(75887)

How long would it take a population growing
at a rate of 3.5% to double itself?
------------
solve (1.035)^x = 2
-----
Take the log to get:
x*log(1.035) = log(2)
--
x = log(2)/log(1.035)
x = 20.148.. years
====
Rounding you would get x = 21 years
===============
Cheers,
Stan H.